OpenStudy (anonymous):

solve the system using any method. x+3y+z=2 4x+y+2z=-4 5x+2y+3z=-2

Directrix (directrix):

Do you have a favorite method you like to use on these?

OpenStudy (anonymous):

no but I tried the addition method but it dindt work out

Directrix (directrix):

Technology not allowed, right?

Directrix (directrix):

@teyo212 Let's go with addition/subtraction. Let's eliminate z. Or, we could go for eliminating the variable you used. Just let me know. @teyo212

OpenStudy (anonymous):

that's fine

OpenStudy (anonymous):

x = -4 y = 0 z = 6

Directrix (directrix):

1) x+3y+z=2 2) 4x+y+2z=-4 3) 5x+2y+3z=-2 I'm numbering these equations as 1, 2, and 3. The object here is to eliminate variable z in all 3 equations. First, we'll eliminate variable z in equations #1 and #2. I'll need you to help because if I just post a solution, I'll be banned from the site. The idea of this site is to have an exchange between the Asker and the Answerer. So, help me out, please. c:

Directrix (directrix):

@kulprit OpenStudy values the Learning process - not the ‘Give you an answer’ process http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

eliminating z from 1) and 2) you arrive at 2x-5y=-8 eliminating z from 1) and 3) you get 2x-7y=-8 ----- (4) clearly y=0 and from (4) you get x=-4 so from (1); z = 2-(-4) = 6 [x y z] = [-4 0 6]

Directrix (directrix):

@teyo212 Would you like to continue?

OpenStudy (anonymous):

yes please

Directrix (directrix):

What do you get when you multiply the first equation by -2? ------ 1) x+3y+z=2 2) 4x+y+2z=-4 3) 5x+2y+3z=-2

OpenStudy (anonymous):

-2x-6y-2z=-4

Directrix (directrix):

Yes. To eliminate z in equations 1 and 2 add what you posted: -2x -6y -2z=-4 4x+ y+ 2z=-4 ============= Post the results of the addition, please. Thanks.

Directrix (directrix):

You are adding your workk to equation 2 to eliminate z. That was the purpose of multiplying the first equation by -2.

OpenStudy (anonymous):

2x-5y=-8

Directrix (directrix):

Correct. We need to keep track of that 2x-5y=-8 *** Now..

Directrix (directrix):

Next up is to eliminate z in equations 1 and 3. To do this, we will multiply equation 1 by -3 and then add that result to equation 3. Your task is to multiply x+3y+z=2 by -3 and to post the result here.

OpenStudy (anonymous):

2x-7y=-4

Directrix (directrix):

-3x - 9y - 3z = -6 5x + 2y + 3z = -2 ============= 2x -7y = -8

Directrix (directrix):

I think you may have forgotten to multiply the 2 of equation 1 by -3.

Directrix (directrix):

Do you agree with this result: 2x -7y = -8 ? ****

OpenStudy (anonymous):

yes

Directrix (directrix):

We have now "reduced" the 3 by 3 system to a 2 by 2 system: 4) 2x-5y=-8 *** 5) 2x -7y = -8 ? **** To eliminate x, multiply equation 4 by -1 and add to equation 5. Please post the result.

Directrix (directrix):

y = ?

OpenStudy (anonymous):

1

Directrix (directrix):

4) 2x-5y=-8 --by -1 --> -2x + 5y = 8 5) 2x -7y = -8 ---------------- -2y = 0 y = 0 ##

Directrix (directrix):

Substitute 0 in for y in this equation and solve for x. 4) 2x-5y=-8 *** @teyo212

OpenStudy (anonymous):

that 2x-7y=-8 suppose to be a -4

Directrix (directrix):

I don't follow what you are saying.

OpenStudy (anonymous):

never mind

Directrix (directrix):

You wrote it earlier in the post: 2x-5y=-8

Directrix (directrix):

Will you do this: Substitute 0 in for y in this equation and solve for x. 4) 2x-5y=-8 ***

OpenStudy (anonymous):

-4

Directrix (directrix):

In equation 1, x+3y+z=2 substitute 0 for y and -4 for x and solve for z. That will be the end of the problem because we will have values for x, y, and z.

OpenStudy (anonymous):

6

Directrix (directrix):

So the solution is (-4, 0 , 6)

OpenStudy (anonymous):

ok thank you for your help

Directrix (directrix):

Glad to help.